An aqueous solution of hydroiodic acid is standardized by titration with a 0.116 M solution of calcium hydroxide.

If 11.2 mL of base are required to neutralize 24.4 mL of the acid, what is the molarity of the hydroiodic acid solution?

M hydroiodic acid
Answer #1

Balanced chemical equation is:

Ca(OH)2 + 2 HI ---> CaI2 + 2 H2O

Here:

M(Ca(OH)2)=0.116 M

V(Ca(OH)2)=11.2 mL

V(HI)=24.4 mL

According to balanced reaction:

2*number of mol of Ca(OH)2 =1*number of mol of HI

2*M(Ca(OH)2)*V(Ca(OH)2) =1*M(HI)*V(HI)

2*0.116*11.2 = 1*M(HI)*24.4

M(HI) = 0.1065 M

Answer: 0.106 M